# Gender-Neutral Large Language Models for Medical Applications: Reducing Bias in PubMed Abstracts

### v1

@@ Abstract
This paper presents a pipeline for mitigating gender bias in large language models (LLMs) used in medical literature by neutralizing gendered occupational pronouns. A dataset of 379,000 PubMed abstracts from 1965-1980 was processed to identify and modify pronouns tied to professions. We developed a BERT-based model, “Modern Occupational Bias Elimination with Refined Training,” or “MOBERT,” trained on these neutralized abstracts, and compared its performance with “1965Bert,” trained on the original dataset. MOBERT achieved a 70% inclusive replacement rate, while 1965Bert reached only 4%. A further analysis of MOBERT revealed that pronoun replacement accuracy correlated with the frequency of occupational terms in the training data. We propose expanding the dataset and refining the pipeline to improve performance and ensure more equitable language modeling in medical applications.
Introduction
Background
@@ Introduction
Despite their potential, though, many LLMs have been shown to contain and perpetuate biases 1, 2, 3, 4, 5. The presence of these biases in LLMs is especially concerning in medical applications, where it can lead to incorrect diagnoses, inappropriate treatment recommendations, and ultimately, unequal healthcare. For example, an LLM fine-tuned on a dataset like PubMed might provide biased diagnostic suggestions if the underlying data contains gendered stereotypes. Gender biases and their effects have already been highlighted in a range of medical practice cases, for topics that include generalized surgical procedures, psychiatry, kidney transplantation, and intensive care treatment, among many other areas 6, 7, 8, 9. Our research focuses specifically on occupational bias in conjunction with gendered pronouns, highlighting the underrepresentation and exclusion of women from traditionally male-dominated professions, a critical area given the concomitant distortion that can result from that in patient care decisions. Numerous instances of this bias are evident in the PubMed dataset, as illustrated in Figure 1. Instead of perpetuating these inequalities, a properly and rigorously trained LLM can mitigate and avoid such dangerous generalizations.
@@
Our method focuses on addressing biases in the training data before the fine-tuning stage of LLM development. By constructing and validating a robust query pipeline that identifies and neutralizes binarily-gendered pronouns linked to occupational terms in medical literature, we aim to create more reliable and fair models. This pipeline includes several key components: a lexicon for identifying gender-specific pronouns, a pronoun resolution query using Llama-3.1, and a classification query to identify occupation-specific subjects. The effectiveness of this approach is demonstrated through the development and evaluation of a BERT-based model, “Modern Occupational Bias Elimination with Refined Training,” or “MOBERT,” trained on gender-neutralized abstracts from PubMed.
Related Work
@@
The issue of bias in word embeddings and LLMs has been widely studied 10, with researchers highlighting how models trained on human-generated corpora often reflect and amplify societal stereotypes 11, 12, as well as proposing both technological and social solutions. Bolukbasi et al. 13 first demonstrated that word embeddings could capture and propagate gender biases, showing that terms like “programmer” were more closely associated with men than women. Subsequent research provided further examples of such biases and explored their widespread implications in the field 14, 15, 16, 17. To counter this effect, Bolukbasi et al. proposed a post-processing technique to debias word embeddings by projecting gender-neutral words into a subspace orthogonal to a gender direction 13. However, their method had limitations, such as requiring a classifier to identify gender-neutral words, which could introduce errors and propagate bias if the classifier itself was flawed.
@@
Building on this, Zhao et al. 18 introduced Gender-Neutral Global Vectors (GN-GloVe), a method that embeds gender information into specific dimensions of word vectors while neutralizing others. This approach improved interpretability and allowed for more effective debiasing by focusing on protected attributes like gender. However, while GN-GloVe effectively reduced direct gender bias, it still left room for improvement in terms of indirect bias and applicability to contextualized word embeddings.
More recently, the focus has shifted to contextualized word embeddings, such as those used in LLMs like BERT and GPT. Studies by Basta et al. 19 and Zhao et al. 20 have explored gender bias in these models, finding that while contextualized embeddings reduce some biases present in static embeddings, they still retain significant levels of bias, especially in how they handle occupations and pronouns in context.
@@
In line with these concerns, recent work has shifted attention toward non-binary and gender-neutral pronouns, as more individuals identify outside the binary gender framework. Although much of the previous research focused on binary gender categories, studies such as Hossain et al. 21 have revealed that large language models struggle significantly with gender-neutral and neo-pronouns, like “they” or “xe.” This highlights the broader issue of representation in training corpora, where non-binary pronouns are often underrepresented, exacerbating the model’s difficulty in handling inclusive language effectively.
@@
Our work builds on these foundations but diverges in its focus on preemptively addressing bias in the training data itself, particularly within the medical domain. By developing a pipeline that identifies and neutralizes gendered language before fine-tuning, we aim to create LLMs that are not only less biased but also more effective in delivering equitable healthcare outcomes. Unlike previous efforts that focused on post-processing or debiasing at the embedding level, our approach integrates bias mitigation into the model training process, addressing both direct and indirect biases more comprehensively.
Methods
Data
@@ Methods
Our data are taken from the MEDLINE 2019 baseline set of PubMed abstracts from 1965 through mid-2018, totaling approximately 29 million abstracts. To locate relevant abstracts for our study, we utilized two lexicons. The first lexicon searched for binarily-gendered pronouns, such as “him,” “her,” and “himself,” ensuring that only complete words were identified. This process reduced the initial set of 29 million abstracts to 687,000 relevant abstract instances.
@@
Annotation Process
Each randomly selected pronoun mention, along with its containing abstract, was uploaded into LabelStudio as part of the broader corpus. The annotation process involved first identifying the noun phrase to which the pronoun referred (defined as “pronoun resolution”) and then classifying that antecedent within the context of the abstract according to the established classification rules. Those rules were set as seen in Table 1.
@@
| Named Individual | Individuals referred to by a proper personal name, which includes capitalized names or redacted names. When an individual is both a patient and a proper name, the “patient” label takes precedence. |
| Occupation | Subjects identified by their profession or job, where they are employed and paid for their work. However, if the individual is also a patient or is named specifically, the “occupation” label is deprioritized, with preference given to labeling them as a patient or by their proper name. |
| Author of the Abstract | The specific individual who wrote the abstract. |
@@
| Other | Any subjects that do not fit into the previous categories. |
In this study, we intentionally avoided using any patient-specific or trial participant-related data to ensure the protection of sensitive information. Consequently, whenever a pronoun refers to a particular patient or group of patients in the abstract, that pronoun was left unchanged. Two annotators (the first author and an intern in the lab) separately labeled each given corpus, then calculated Cohen’s Kappa, a measure quantifying the level of agreement before reconciliation.
Example annotations from the corpus are shown in Figure 1. Descriptive statistics of the annotated corpora are provided in Table 2.
| Pronoun- and | ||||
| Pronoun-Only Corpus | Lexicon-Derived Corpus | |||
| Number of Abstracts | 250 | Number of Abstracts | 500 | |
| Patient/Trial Participant | 28 | Patient/Trial Participant | 323 | |
| Named Individual | 62 | Named Individual | 115 | |
| Occupation | 97 | Occupation | 24 | |
| Author of the Abstract | 56 | Author of the Abstract | 19 | |
| Animal | 0 | Animal | 7 | |
| Other | 7 | Other | 12 | |
Pronoun Resolution Query
@@
Our pipeline initially employed a Llama-3.1-405b query to extract the subject associated with each pronoun (pronoun resolution). A randomly selected corpus of 500 pronoun instances was chosen from the relevant abstracts and each pronoun’s respective antecedent was located and double-annotated with a Cohen’s Kappa of 0.9000. Selected examples and the overall makeup of this corpus can be seen in both Figure 1 and Table 2. The Llama-3.1 query was then run on the non-annotated corpus and the pronoun instances and their respective antecedent query output were cross-referenced with the ground-truth annotations. The numerical outcome of this process is shown in Table 3.
| Initial 500 abstracts | Accuracy |
| 0.9881 |
Lexicon Validation
Once the accuracy of the pronoun resolution query is verified, a lexicon must be designed to identify occupational antecedents with greater frequency. The fidelity of the lexicon must also be evaluated in an iterative manner, through applying it to search for occupational terms within 500 antecedents identified by the pronoun resolution query, the process for which is described below.
Based on observations made in the pronoun resolution query stage, it was found that a lexicon was needed that was specifically designed to identify occupational antecedents with greater frequency. In order to ensure the broad applicability of this lexicon to the relevant sample set, it was initially derived from the synset relations of “professional” in WordNet and then refined to include case-sensitive acronyms (e.g., “rn” versus “RN”). To validate the lexicon for this use, we applied it to search for occupational terms within the 500 antecedents identified by the pronoun resolution query. The filtered results were then compared to the “occupation” labels in the ground-truth annotations. In our case, the application achieved a perfect recall score of 1.0000, confirming that the lexicon reliably identifies occupational antecedents. This validation supports its further use in refining and testing the pipeline on relevant data.
Classification Query
@@
Using the validated lexicon described above, the results of our antecedent query can be successfully filtered. In this application, 250 pronoun instances were extracted, along with their corresponding antecedents, that included occupational terms. These 250 instances and the text of the abstract in which they appeared must also be examined and tested for accuracy. Those instances were double-annotated and reconciled with a Cohen’s Kappa of 0.9470. After the annotations were completed, an additional Llama-3.1-405b query must then be employed to similarly label each antecedent in the context of its abstract. Precision, recall, and the F1 score were calculated between the generated labels and the ground-truth labels. The numerical outcome of this process is shown in Table 4 and an example of an antecedent versus classification query is given in Table 5. The primary focus of the classification query is identifying instances of occupations. The other categories were annotated to provide context for pronoun usage distribution, and offer potential for deeper analysis in future work.
@@
| Macro Weighted Average | 0.943 | 0.932 | 0.9349 |
| Antecedent Query | System Content | User Content |
| You are a helpful assistant with identifying the direct antecedent of a pronoun. Here is your antecedent_background knowledge: {background}.* | Identify the direct antecedent of the pronoun marked with [START] and [END] in the following abstract: {highlighted_abstract}. Only answer with the antecedent. | |
| Classification Query | System Content | User Content |
| {antecedent} is the output of Antecedent Query | You are a helpful assistant following these classification rules {rules}.** | In the following abstract, classify which label the noun “{antecedent}” in the context of the abstract {highlighted_abstract} is referring to: “patient,” “occupation,” “named individual,” “author,” “animal,” or “other.” Only output the label, no other text. |
*This background information consists of antecedent grammatical rules established by 22.
@@
Results: Application of the Pipeline 1965-1980
Pronoun Neutralization Case Study
@@
We tested the effect of our pipeline on a corpus of the 379,000 PubMed abstracts from 1965-1980, hypothesizing that these texts would show a greater prevalence of singular gendered pronouns, based on a qualitative examination of a random sample set of the abstracts. After processing this corpus through our pipeline, the output consisted of all of the abstracts containing singular gendered pronouns referring to a labeled occupational antecedent, along with the pronoun character offsets. These were the pronouns deemed acceptable to neutralize as they are solely linked to an occupational term, and the abstract will not lose any medically necessary context (e.g., the gender of a trial participant).
To neutralize these pronouns, NLTK resources were downloaded and used for text tokenization and part-of-speech tagging. A pronoun-mapping dictionary was defined to convert gender-specific pronouns to gender-inclusive equivalents, with functions set up to identify, classify, and replace pronouns within the text. This dictionary also accounted for compound pronouns, such as “he or she” or “his/her,” and grouped them for replacement accordingly. Gender-specific pronouns (both compounded and non-compounded) at the given offsets were replaced with the gender-inclusive counterparts “they/them/theirs.” Examples of phrases that would or would not be tapped for replacement, and the resulting modifications, are shown in Table 6.
@@
To determine the success of this replacement, we trained two separate base uncased BERT models. The first model, named 1965Bert, was trained on the original, unmodified dataset of the 379,000 PubMed abstracts from 1965-1980. The second model, denoted “Modern Occupational Bias Elimination with Refined Training,” or “MOBERT,” was trained on a similar dataset of the 379,000 abstracts, but with 1,400 abstracts identified and modified with the newly introduced gender-neutral tokens. Additionally, these 1,400 abstracts were analyzed to identify the most frequently-occurring occupational terms in relevant antecedents. The top five terms identified were “physician,” “surgeon,” “doctor,” “practitioner,” and “nurse.” Both models were trained for 3 epochs with a batch size of 4 per device, using a mixed precision (fp16) configuration across multiple GPUs. Training logs were saved at regular intervals, with models checkpointed every 10,000 steps.
@@
To further assess the models, we conducted a masked language modeling test using 50 sentences from our initial annotated corpus of 500 abstracts, ensuring that each randomly selected sentence contained gendered pronouns from post-1980 texts. Importantly, the models were not trained on the data used in these tasks, ensuring an independent evaluation of their performance. The testing corpus was assembled by selecting ten sentences for each of the five most frequent occupational terms identified, resulting in 50 sentences. In each sentence, a [MASK] token was inserted in place of a pronoun, and the model was tasked with predicting the correct pronoun when given respective options of he/him/his, she/her/hers, and they/them/theirs.
@@
Outcomes
We compared the results of this masking test between Bert-Base (the original untrained model), and microsoft/Biomed- NLP-PubMedBERT-base-uncased-abstract, 1965Bert, and MOBERT (all three of which are trained upon Bert-Base with their respective training data). Examples of this masking test and the corresponding outcomes can be seen in Table 7, with overall results shown in Table 8. Percentages are to indicate the share of sentences replaced with a gender inclusive pronoun of they/them/theirs. For example, if Bert-Base replaces 40% of masked pronoun instances with a gender inclusive pronoun, 1965Bert replaces 4% of those same instances with a gender inclusive pronoun.

### v2

@@ Abstract
This paper presents a pipeline for mitigating gender bias in large language models (LLMs) used in medical literature by neutralizing gendered occupational pronouns. A set of 379,000 PubMed abstracts from 1965-1980 was processed to identify and modify pronouns tied to professions. We developed a BERT-based model, “Modern Occupational Bias Elimination with Refined Training,” or “MOBERT,” trained on these neutralized abstracts, and compared it with “1965BERT,” trained on the original dataset. MOBERT achieved a 70% inclusive replacement rate, while 1965BERT reached only 4%. A further analysis of MOBERT revealed that pronoun replacement accuracy correlated with the frequency of occupational terms in the training data. We propose expanding the dataset and refining the pipeline to improve performance and ensure more equitable language modeling in medical applications.
Gender-Neutral Large Language Models for Medical Applications: Reducing Bias in PubMed Abstracts
Elizabeth Schaefer Yale University Department of Computer Science New Haven, CT Elizabeth.Schaefer@yale.edu Kirk Roberts UTHealth Houston McWilliams School of Biomedical Informatics, Houston, TX Kirk.Roberts@uth.tmc.edu
1 Introduction
1.1 Background
@@
Despite their potential, though, many LLMs have been shown to contain and perpetuate biases Kotek et al. (2023); Liu et al. (2022); Abid et al. (2021); Prakash and Lee (2023); Bai et al. (2024); Zack et al. (2024); Bedi et al. (2024); Degelin (2024). The presence of these biases in LLMs is especially concerning in medical applications, where it can lead to incorrect diagnoses, inappropriate treatment recommendations, and ultimately, unequal healthcare. For example, an LLM fine-tuned on a dataset like PubMed might provide biased diagnostic suggestions if the underlying data contain gendered stereotypes. Gender biases and their effects have already been highlighted in a range of medical practice cases, for topics that include generalized surgical procedures, psychiatry, kidney transplantation, and intensive care treatment, among many other areas Ruiz and Verbrugge (1997); van Daal et al. (2020); Lim et al. (2021); Merdji et al. (2023); Omar et al. (2024). Our research focuses specifically on occupational bias in conjunction with gendered pronouns, highlighting the underrepresentation and exclusion of women from traditionally male-dominated professions, a critical area given the concomitant distortion that can result from that in patient care decisions. Numerous instances of this bias are evident in the PubMed dataset, as illustrated in Figure 1. Instead of perpetuating these inequalities, a properly and rigorously trained LLM can mitigate and avoid such dangerous generalizations.
@@
Our method focuses on addressing biases in the training data before the fine-tuning stage of LLM development. By constructing and validating a robust query pipeline that identifies and neutralizes binarily-gendered pronouns linked to occupational terms in medical literature, we aim to create more reliable and fair models. This aligns with concerns raised by Bender et al. (2021), who emphasize that training on biased corpora can amplify existing societal stereotypes in unintended ways. By modifying the dataset at the source rather than post-processing model outputs, we directly address these concerns and create a more stable foundation for fairness in medical NLP. Our pipeline includes several key components: a lexicon for identifying gender-specific pronouns, a pronoun resolution query, and a classification query to identify occupation-specific subjects. Both queries were conducted using Meta’s Llama-3.1-405b model Meta AI (2024), which we elected to use because it was the most advanced Llama model available at the time, and offered improved reasoning and accuracy for the task compared to previous models. The effectiveness of this approach is demonstrated through the development and evaluation of a BERT-based model, “Modern Occupational Bias Elimination with Refined Training,” or “MOBERT,” trained on gender-neutralized abstracts from PubMed.
1.2 Related Work
@@
The issue of bias in word embeddings and LLMs has been widely studied Pessach and Shmueli (2022), with researchers highlighting how models trained on human-generated corpora often reflect and amplify societal stereotypes Dev et al. (2023); Ungless et al. (2022), as well as proposing both technological and social solutions. Bolukbasi et al. (2016) first demonstrated that word embeddings could capture and propagate gender biases, showing that terms like “programmer” were more closely associated with men than women. Subsequent research provided further examples of such biases and explored their widespread implications in the field Ray (2023); Bommasani et al. (2021); Mehrabi et al. (2021); An et al. (2024); Pervez and Titus (2024). To counter this effect, Bolukbasi et al. (2016) proposed a post-processing technique to debias word embeddings by projecting gender-neutral words into a subspace orthogonal to a gender direction. However, their method had limitations, such as requiring a classifier to identify gender-neutral words, which could introduce errors and propagate bias if the classifier itself was flawed.
@@
Zhao et al. (2018) subsequently introduced GN-GloVe, a method that embeds gender information into specific dimensions of word vectors while neutralizing others. This approach improved interpretability and allowed for more effective debiasing by focusing on protected attributes like gender. However, while GN-GloVe effectively reduced direct gender bias, it left room for improvement in terms of indirect bias and applicability to contextualized word embeddings. As Blodgett et al. (2020) highlight, many prior bias-mitigation efforts rely on lexicon-based heuristics or simple vector-space transformations, which may fail to generalize to real-world applications. Our work moves beyond this by integrating pronoun resolution and context-aware classification before model fine-tuning.
@@
More recently, the focus has shifted to contextualized word embeddings, such as those used in LLMs like BERT and GPT. Basta et al. (2019) and Zhao et al. (2019) explored gender bias in these models, finding that while contextualized embeddings reduce some biases present in static embeddings, they still retain significant levels of bias, especially in how they handle occupations and pronouns in context. In line with these concerns, recent work has shifted attention toward non-binary and gender-neutral pronouns, as more individuals identify outside the binary gender framework. Although much of the previous research focused on binary gender categories, studies such as Hossain et al. (2023) have revealed that large language models struggle significantly with gender-neutral and neo-pronouns, like “they” or “xe.” This highlights the broader issue of representation in training corpora, where non-binary pronouns are often underrepresented, exacerbating the model’s difficulty in handling inclusive language effectively.
Beyond post-processing and embedding-based debiasing methods, recent research has explored direct modifications to training data as a strategy for mitigating gender bias in biomedical NLP. Agmon et al. (2024) introduced TeDi-BERT, a model that applies temporal distribution matching to adjust how gender is represented in historical clinical trial abstracts, ensuring that language models trained on older data align more closely with contemporary gender distributions. Their approach highlights the importance of modifying training corpora before model training to prevent biased language from propagating in downstream applications.
@@
Our work builds on these foundations but diverges in its focus on occupational bias in medical literature. Rather than aligning embeddings across different time periods, we develop a pipeline that systematically identifies and neutralizes gendered occupational pronouns before model fine-tuning. Through this methodology, we aim to create LLMs that are not only less biased but also more effective in delivering equitable healthcare outcomes. Unlike previous efforts that focused on post-processing or debiasing at the embedding level, our approach integrates bias mitigation into the model training process, addressing both direct and indirect biases more comprehensively.
| Dataset A: Pronoun-only corpus | Dataset B: Pronoun- and Lexicon-Derived Corpus | |
| Total Number of Abstracts | 250 | 500 |
| Patient/Trial Participant | 28 | 323 |
| Named Individual | 62 | 115 |
| Occupation | 97 | 24 |
| Author of the Abstract | 56 | 19 |
| Animal | 0 | 7 |
| Other | 7 | 12 |
2 Methods
2.1 Data
@@
Our data are taken from the MEDLINE 2019 baseline set of PubMed abstracts from 1965 through mid-2018, totaling approximately 29 million abstracts. We utilized two lexicons to locate relevant abstracts for our study. The first lexicon searched for binarily-gendered pronouns, such as “him,” “her,” and “himself,” ensuring that only complete words were identified. This process reduced the initial set of 29 million abstracts to 687,000 relevant abstract instances. A second lexicon, designed to identify occupational terms, was applied only for testing purposes, allowing us to evaluate classification performance across a broader range of occupation-related pronouns. This second lexicon was not used in the case study dataset and did not affect the pronoun-neutralization process. For clarity, we designate the “Pronoun-only corpus” as Dataset A and the “Pronoun- and Lexicon-Derived Corpus” as Dataset B. These names will be used throughout the paper when referring to the annotated datasets, as in Table 1.
@@
| Antecedent category | Category definition |
@@
| Named Individual | Individuals referred to by a proper personal name, which includes capitalized names or redacted names. |
| Occupation | Individuals, real or abstract, identified by their profession or job, where they are employed and paid for their work. |
| Author of the Abstract | An author of the paper. |
@@
| Other | Any instance that does not fit into the previous categories. |
2.2 Annotation Process
After constructing Datasets A and B, we proceeded with a two-step annotation process involving pronoun resolution and antecedent classification. This annotation process (utilizing LabelStudio Tkachenko and contributors (2020)) involved first identifying the noun phrase to which the pronoun referred (defined as “pronoun resolution”) and then classifying that antecedent within the context of the abstract according to the established classification rules. Those rules were set as seen in Table 2. In this study, we intentionally avoided neutralizing pronouns when referring to patients or trial participants, as well as in contexts where biological sex is medically relevant. Certain conditions, such as prostate or ovarian cancer, are inherently gendered, and so de-gendering such references could hamper a model’s medical reasoning. Consequently, whenever a pronoun refers to a particular patient or group of patients in the abstract, that pronoun was left unchanged. Two annotators (the first author and an intern in the lab) separately labeled each given corpus, then calculated Cohen’s Kappa, a measure quantifying the level of agreement before reconciliation. Example annotations from the corpus are shown in Figure 1. Descriptive statistics of the annotated corpora from Dataset A and Dataset B are provided in Table 1.
2.3 Pronoun Resolution Query
@@
As the first step in our automated pipeline, we used a Llama-3.1-405b query to determine the subject associated with each pronoun, a process known as pronoun resolution. This step involved determining the noun or noun phrase to which a given pronoun referred within the abstract’s context, ensuring that additional descriptive clauses were excluded. The full structured prompt used in this query is detailed in Table 4. To evaluate this prompt-based model, a randomly selected corpus of 500 pronoun instances was chosen from the relevant abstracts and each pronoun’s respective antecedent was located and double-annotated with a Cohen’s Kappa of 0.9000. Selected examples and the overall makeup of this corpus can be seen in both Figure 1 and Table 2.
2.4 Lexicon Validation
Following pronoun resolution, it proved useful to define a mechanism to distinguish occupational antecedents from other noun phrases. While the LLM was highly effective in identifying pronoun antecedents, we found that only a small percentage of gendered pronouns were actually attributable to professions, the core focus of our task. This data imbalance made it difficult to obtain a sufficient sample of occupation-related pronoun instances for evaluation. To address this, we developed a lexicon specifically designed to increase the frequency of identified occupational antecedents. The lexicon was initially derived from the synset relations of “professional” in WordNet, incorporating common occupational terms and case-sensitive acronyms (e.g., “rn” vs. “RN”). To validate the lexicon’s efficacy, it was applied to the 500 resolved antecedents, filtering for occupational terms.
2.5 Classification Query
@@
Using the validated lexicon described above, the results of our antecedent query can be successfully filtered. In this application, 250 pronoun instances were extracted from the data (primarily from Dataset B with the applied lexicon), along with their corresponding antecedents, that included occupational terms. These 250 instances and the text of the abstract in which they appeared must also be examined and tested for accuracy. Those instances were double-annotated and reconciled with a Cohen’s Kappa of 0.9470. After annotator reconciliation, we used the Llama-3.1-405b model to classify each antecedent according to the same labeling rules, enabling comparison between human and model performance.
@@
| Macro Weighted Avg. | 0.943 | 0.932 | 0.9349 |
| System Content | User Content | |
| Pronoun Resolution Query | You are a helpful assistant with identifying the direct antecedent of a pronoun. Here is your antecedent_background knowledge: {background}’*. | Identify the direct antecedent of the pronoun marked with [START] and [END] in the following abstract: {highlighted_abstract}. Only answer with the antecedent. |
| Classification Query (where antecedent is the output of Antecedent Query) | You are a helpful assistant following these classification rules {rules}.** | In the following abstract, classify which label the noun “{antecedent}” in the context of the abstract {highlighted_abstract} is referring to: “patient,” “occupation,” “named individual,” “author,” “animal,” or “other.” Only output the label, no other text. |
*This background information consists of antecedent grammatical rules established by Fordham Fordham University (2024).
@@
2.6 Pronoun Neutralization Process
After identifying gendered pronouns linked to occupational terms, the next step was neutralization. NLTK was used for tokenization and part-of-speech tagging. A pronoun-mapping dictionary was developed to replace gendered pronouns with their gender-neutral counterparts, such as ‘they/them/theirs.’ This dictionary accounted for compound pronouns (e.g., he or she they) and handled replacements while preserving sentence structure. Pronouns flagged for neutralization were modified only when they referred to occupational antecedents, ensuring no changes were made to pronouns referring to patients or trial participants. This distinction was critical for maintaining medically relevant context in abstracts where sex-specific conditions (e.g., prostate cancer) were discussed. Examples of pronoun replacements and contexts that were preserved are presented in Table 5. The following section presents the results obtained from applying the three-stage pipeline on our annotated datasets, showcasing the effectiveness of our approach in neutralizing gendered pronouns.
@@ 2.6 Pronoun Neutralization Process
3 Results
3.1 Pipeline Performance Metrics
To evaluate the performance of our pipeline, we analyzed two separate annotated datasets. Dataset A (Pronoun-only corpus) was used to assess the pronoun resolution component, while Dataset B (Pronoun- and Lexicon-Derived corpus) was utilized for the lexicon validation and classification queries. First, using Dataset A, we applied our pronoun resolution query on the non-annotated corpus. The Llama-3.1-405b query was run and the resulting pronoun instances and their corresponding antecedent outputs were cross-referenced with the ground-truth annotations. This comparison yielded an accuracy of 0.9881 on the initial 500 abstracts, demonstrating that our pronoun resolution method reliably identifies antecedents. Next, with Dataset B, we validated our lexicon for identifying occupational terms by applying it to the 500 antecedents obtained from the pronoun resolution query. The filtered results were then compared with the ‘occupation’ labels in the ground-truth annotations, achieving a perfect recall score of 1.0000. This confirms that our lexicon effectively identifies occupational antecedents for the classification task. Finally, still using Dataset B, we assessed the performance of our Llama-3.1-405b classification query by calculating precision, recall, and the F1 score between the generated labels and the ground-truth labels. The numerical outcomes of this process are presented in Table 3, and an example of an antecedent versus classification query is provided in Table 4. These results confirm the high performance of our classification approach in accurately distinguishing occupational pronoun instances from other categories. Together, these performance metrics validate the robustness of our pipeline, linking each methodological step to successful outcomes in resolving, filtering, and classifying gendered pronoun instances.
3.2 Pronoun Neutralization Case Study
@@
We tested the effect of our pipeline on a corpus of the 379,000 PubMed abstracts from 1965-1980, hypothesizing that these texts would show a greater prevalence of singular gendered pronouns, based on a qualitative examination of a random sample set of the abstracts. After processing this corpus, pronouns linked to occupational antecedents were neutralized in 1,400 abstracts.
@@
To determine the success of this replacement, we trained two separate base uncased BERT models Devlin et al. (2019). The first model, named 1965BERT, was trained on the original, unmodified dataset of the 379,000 PubMed abstracts from 1965-1980. The second model, denoted “Modern Occupational Bias Elimination with Refined Training,” or MOBERT, was trained on a similar dataset of the 379,000 abstracts, but with 1,400 abstracts identified and modified with the newly introduced gender-neutral tokens. Additionally, these 1,400 abstracts were analyzed to identify the most frequently-occurring occupational terms in relevant antecedents. The top five terms identified were “physician,” “surgeon,” “doctor,” “practitioner,” and “nurse.” Both models were trained for three epochs with a batch size of four per device, using a mixed precision (fp16) configuration across multiple graphics processing units. Training logs were saved at regular intervals, with models checkpointed every 10,000 steps.
@@
To further assess the models, we conducted a masked language modeling test using 50 sentences from our initial annotated corpus of 500 abstracts, ensuring that each randomly selected sentence contained gendered pronouns from post-1980 texts. Importantly, the models were not trained on the data used in these tasks, ensuring an independent evaluation of their performance. The testing corpus was assembled by selecting ten sentences for each of the five most frequent occupational terms identified, resulting in 50 sentences. In each sentence, a [MASK] token was inserted in place of a pronoun, and the model was tasked with predicting the correct pronoun when given respective options of ‘he/him/his,’ ‘she/her/hers,’ and ‘they/them/theirs.’
@@
3.3 Outcomes
We compared the results of this masking test between BERT-Base (the untrained model), PubMedBERT, 1965BERT, and MOBERT (all three of which are trained upon BERT-Base with their respective training data) Gu et al. (2020). Examples of this masking test and the corresponding outcomes are in Table 6, with overall results shown in Table 7. Percentages indicate the proportion of sentences in which gender-inclusive pronouns (‘they/them/theirs’) replaced gendered pronouns. For example, if BERT-Base replaces 40% of masked pronoun instances with a gender inclusive pronoun, 1965BERT replaces 4% of those same instances with a gender inclusive pronoun. The MOBERT results were further analyzed to determine a relationship between the frequency of the occupational term in the training data and the accuracy of replacement, as shown in Table 8.
4 Discussion
4.1 Principal Results
The application of our gender-neutralization pipeline to the 1965-1980 PubMed abstracts has demonstrated its potential to significantly reduce occupational gender bias in large language models. By introducing gender-neutral pronouns reconciled with occupational terms in 1,400 abstracts, we successfully trained a model, MOBERT, that demonstrated a 70% success rate in predicting inclusive pronouns in a masked language modeling task. This result far exceeds the 4% success rate of 1965BERT, a model trained on unmodified texts from the same period, and highlights the importance of correcting biased data at the training stage. MOBERT’s performance also surpassed that of both the base model, BERT-Base, which exhibited a 40% success rate, and PubMedBERT, a model trained on the complete PubMed dataset without gender-neutralization, which achieved only a 20% inclusive successive rate. These comparisons underscore the critical role of targeted intervention in mitigating bias in language models.